A franchise of automotive technicians has a contract arrangement for a rewards program with a car manufacturing company called BMC. For every $4.00 of services performed by an automotive technician at anyone of the franchise locations, the customer is awarded Car Dollars. The Car Dollars can be redeemed towards the purchase of a motor vehicle manufactured by BMC. In total the automotive technicians performed $20,200,000 worth of services in 2020. BMC have estimated the stand-alone selling price of $0.35 and for each Car Dollar awarded, the automotive technicians will pay $0.30 to the BMC for each Car Dollar awarded to its customers.

Collapse question part

(a)

Determine how the $20,200,000 transaction price would be allocated to the two performance obligations (the automotive services and the Car Dollars earned by the customers). (Round percentage of total selling price calculations to 1 decimal place, e.g. 15.2% and final answers to 0 decimal places, e.g. 125.)
Automotive Services $

Car Dollars $
Answer #1

No. of car dollars awarded =  $20,200,000/$4 = 5,050,000

Standalone selling price = 5050000*0.35 = 1767500

%

Automotive Services $20200000 92.0%

Car Dollars $ 1767500 8.0%

Total $21967500 100%

Total revenue would be divided into stand alone selling prices.

So, revenue to be recognised for Automative services = 20200000*92% = 18574712.6

And for Car dollars = 20200000*8% = 1625287.4